OpenStudy (anonymous):

a culture of bacteria contains 10000 bacteria initially. after an house the bacteria count is 25000. find the doubling time to the nearest minute.

OpenStudy (anonymous):

i am going to make a guess that "house" is an hour

OpenStudy (anonymous):

yes oops!

OpenStudy (anonymous):

ok we can do this a couple ways, but the snap way is this

OpenStudy (anonymous):

\[25000\div10000=2.5\] set \[(2.5)^{\frac{t}{60}}=2\] and solve for \(t\)

OpenStudy (anonymous):

i put a 60 in the denominator to convert the units from hours to minutes at the beginning. we could have converted at the end no matter solve via \[\frac{t}{60}=\frac{\ln(2)}{\ln(2.5)}\] and so \[t=\frac{60\times \ln(2)}{\ln(2.5)}\] and a calculator

OpenStudy (anonymous):

45.39

OpenStudy (anonymous):

you can also model this as \[1000e^{rt}\] and then you have to find \(r\) and then the doubling time, but it seems like a large waste of time to me, since the numbers i used in the expression \[(2.5)^{\frac{t}{60}}\] i got from reading, not computing

OpenStudy (anonymous):

seems like a reasonable estimate to me

OpenStudy (anonymous):

so then to figure out how many bacteria would be there after 3 hours, you would do what?

OpenStudy (anonymous):

the model is \[1000\times (2.5)^{\frac{t}{60}}\]

OpenStudy (anonymous):

you mean 10000 right

OpenStudy (anonymous):

for 3 hours replace \(t\) by 180 or if you are working in hours compute \[10000\times (2.5)^3\]

OpenStudy (anonymous):

oh yeah i guess it is 10000 i missed a zero

OpenStudy (anonymous):

156,250

OpenStudy (anonymous):

hope it is clear how i arrived at \((2.5)^{\frac{t}{60}}\) almost instantly the population increased by a factor of \(2.5\) in 60 minutes

OpenStudy (anonymous):

yeah it is thank you

OpenStudy (anonymous):

that answer cannot be right

OpenStudy (anonymous):

oh maybe it is, let me check

OpenStudy (anonymous):

yes, it is right, sorry

OpenStudy (anonymous):

ok